\begin{thebibliography}{10}

\bibitem{brazdil2008metalearning}
Pavel Brazdil, Christophe~Giraud Carrier, Carlos Soares, and Ricardo Vilalta.
\newblock {\em Metalearning: Applications to data mining}.
\newblock Springer Science \& Business Media, 2008.

\bibitem{Fonseca2014Evaluation}
J.~Fonseca, M.~Vieira, and H.~Madeira.
\newblock Evaluation of web security mechanisms using vulnerability \& attack
  injection.
\newblock {\em Dependable \& Secure Computing IEEE Transactions on},
  11(5):440--453, 2014.

\bibitem{hydara2015current}
Isatou Hydara, Abu Bakar~Md Sultan, Hazura Zulzalil, and Novia Admodisastro.
\newblock Current state of research on cross-site scripting (xss)--a systematic
  literature review.
\newblock {\em Information and Software Technology}, 58:170--186, 2015.

\bibitem{WAAR2014}
Imperva.
\newblock Web application attack report 2014.
\newblock
  \url{https://www.imperva.com/docs/HII_Web_Application_Attack_Report_Ed6.pdf}.

\bibitem{WAAR2015}
Imperva.
\newblock Web application attack report 2015.
\newblock
  \url{https://www.imperva.com/docs/HII_Web_Application_Attack_Report_Ed6.pdf}.

\bibitem{WAAR2016}
Imperva.
\newblock Web application attack report 2016.
\newblock
  \url{https://www.imperva.com/docs/HII_Web_Application_Attack_Report_Ed6.pdf}.

\bibitem{Juvonen2015Online}
Antti Juvonen, Tuomo Sipola, and Timo Inen.
\newblock Online anomaly detection using dimensionality reduction techniques
  for http log analysis.
\newblock {\em Computer Networks the International Journal of Computer \&
  Telecommunications Networking}, 91(C):46--56, 2015.

\bibitem{Kozakevicius2015URL}
Alice Kozakevicius, Cristian Cappo, Bruno~A. Mozzaquatro, Raul~Ceretta Nunes,
  and Christian~E. Schaerer.
\newblock Url query string anomaly sensor designed with the bidimensional haar
  wavelet transform.
\newblock {\em International Journal of Information Security}, 14(6):561--581,
  2015.

\bibitem{Kruegel2003Anomaly}
Christopher Kruegel and Giovanni Vigna.
\newblock Anomaly detection of web-based attacks.
\newblock In {\em ACM Conference on Computer and Communications Security},
  pages 251 -- 261, 2003.

\bibitem{Kruegel2005A}
Christopher Kruegel, Giovanni Vigna, and William Robertson.
\newblock A multi-model approach to the detection of web-based attacks.
\newblock {\em Computer Networks the International Journal of Computer \&
  Telecommunications Networking}, 48(5):717--738, 2005.

\bibitem{lawal2016systematic}
MA~Lawal, Abu Bakar~Md Sultan, and Ayanloye~O Shakiru.
\newblock Systematic literature review on sql injection attack.
\newblock {\em International Journal of Soft Computing}, 11(1):26--35, 2016.

\bibitem{Meng2014Adaptive}
Yuxin Meng and Lam~For Kwok.
\newblock Adaptive blacklist-based packet filter with a statistic-based
  approach in network intrusion detection.
\newblock {\em Journal of Network \& Computer Applications}, 39(1):83--92,
  2014.

\bibitem{Prokhorenko2016Web}
Victor Prokhorenko, Kim Kwang~Raymond Choo, and Helen Ashman.
\newblock Web application protection techniques: A taxonomy.
\newblock {\em Journal of Network \& Computer Applications}, 60:95--112, 2016.

\bibitem{Robertson2006Using}
William~K. Robertson, Giovanni Vigna, Christopher Kruegel, and Richard~A.
  Kemmerer.
\newblock Using generalization and characterization techniques in the
  anomaly-based detection of web attacks.
\newblock In {\em Network and Distributed System Security Symposium, NDSS 2006,
  San Diego, California, Usa}, 2006.

\bibitem{Song2009Spectrogram}
Yingbo Song, Angelos~D. Keromytis, and Salvatore~J. Stolfo.
\newblock Spectrogram: A mixture-of-markov-chains model for anomaly detection
  in web traffic.
\newblock In {\em Network and Distributed System Security Symposium, NDSS 2009,
  San Diego, California, Usa, February - February}, pages 121--135, 2009.

\bibitem{ISTR2016}
Symantec.
\newblock Internet security threat report 2016.
\newblock
  \url{https://www.symantec.com/content/dam/symantec/docs/reports/istr-21-2016-en.pdf}.

\bibitem{Ghost}
Symantec.
\newblock Team ghostshell hacking group back with a bang.
\newblock
  \url{https://www.symantec.com/connect/blogs/team-ghostshell-hacking-group-back-bang}.

\bibitem{Vapnik2006Estimation}
Vladimir Vapnik.
\newblock {\em Estimation of Dependences Based on Empirical Data}.
\newblock Springer New York, 2006.

\bibitem{Wang2014Autonomic}
Wei Wang, Thomas Guyet, René Quiniou, Marie~Odile Cordier, Florent Masseglia,
  and Xiangliang Zhang.
\newblock Autonomic intrusion detection: Adaptively detecting anomalies over
  unlabeled audit data streams in computer networks.
\newblock {\em Knowledge-Based Systems}, 70:103--117, 2014.

\bibitem{Zhang2015Detecting}
Ji~Zhang, Hongzhou Li, Qigang Gao, Hai Wang, and Yonglong Luo.
\newblock Detecting anomalies from big network traffic data using an adaptive
  detection approach.
\newblock {\em Information Sciences An International Journal}, 318(C):91--110,
  2015.

\end{thebibliography}


\begin{thebibliography}{10}

\bibitem{Ariu2011HMMPayl}
Davide Ariu, Roberto Tronci, and Giorgio Giacinto.
\newblock Hmmpayl : An intrusion detection system based on hidden markov
  models.
\newblock {\em Computers \& Security}, 30(4):221--241, 2011.

\bibitem{Dahse2014Static}
Johannes Dahse and Thorsten Holz.
\newblock Static detection of second-order vulnerabilities in web applications.
\newblock 2014.

\bibitem{Garcia2015Automatic}
P.~Garcia-Teodoro, J.~E. Diaz-Verdejo, J.~E. Tapiador, and
  R.~Salazar-Hernandez.
\newblock Automatic generation of http intrusion signatures by selective
  identification of anomalies.
\newblock {\em Computers \& Security}, 55(C):159--174, 2015.

\bibitem{Howaed2014psigene}
Gaspar~Modelo Howard, Christopher~N Gutierrez, Fahad~A Arshad, Saurabh Bagchi,
  and Yuan Qi.
\newblock psigene: Webcrawling to generalize sql injection signatures.
\newblock In {\em 2014 44th Annual IEEE/IFIP International Conference on
  Dependable Systems and Networks}, pages 45--56. IEEE, 2014.

\bibitem{WAAR2014}
Imperva.
\newblock Web application attack report 2014.
\newblock
  \url{https://www.imperva.com/docs/HII_Web_Application_Attack_Report_Ed6.pdf}.

\bibitem{WAAR2015}
Imperva.
\newblock Web application attack report 2015.
\newblock
  \url{https://www.imperva.com/docs/HII_Web_Application_Attack_Report_Ed6.pdf}.

\bibitem{WAAR2016}
Imperva.
\newblock Web application attack report 2016.
\newblock
  \url{https://www.imperva.com/docs/HII_Web_Application_Attack_Report_Ed6.pdf}.

\bibitem{Jamdagni2013Review}
Aruna Jamdagni, Zhiyuan Tan, Xiangjian He, Priyadarsi Nanda, and Ren~Ping Liu.
\newblock Repids: A multi tier real-time payload-based intrusion detection
  system.
\newblock {\em Computer Networks}, 57(3):811--824, 2013.

\bibitem{Juvonen2015Online}
Antti Juvonen, Tuomo Sipola, and Timo Inen.
\newblock Online anomaly detection using dimensionality reduction techniques
  for http log analysis.
\newblock {\em Computer Networks the International Journal of Computer \&
  Telecommunications Networking}, 91(C):46--56, 2015.

\bibitem{Kozakevicius2015URL}
Alice Kozakevicius, Cristian Cappo, Bruno~A. Mozzaquatro, Raul~Ceretta Nunes,
  and Christian~E. Schaerer.
\newblock Url query string anomaly sensor designed with the bidimensional haar
  wavelet transform.
\newblock {\em International Journal of Information Security}, 14(6):561--581,
  2015.

\bibitem{Kruegel2003Anomaly}
Christopher Kruegel and Giovanni Vigna.
\newblock Anomaly detection of web-based attacks.
\newblock In {\em ACM Conference on Computer and Communications Security},
  pages 251 -- 261, 2003.

\bibitem{Kruegel2005A}
Christopher Kruegel, Giovanni Vigna, and William Robertson.
\newblock A multi-model approach to the detection of web-based attacks.
\newblock {\em Computer Networks the International Journal of Computer \&
  Telecommunications Networking}, 48(5):717--738, 2005.

\bibitem{Livshits2010Towards}
Benjamin Livshits and Stephen Chong.
\newblock Towards fully automatic placement of security sanitizers and
  declassifiers.
\newblock {\em Acm Sigplan Notices}, 48(48):385--398, 2010.

\bibitem{Meng2014Adaptive}
Yuxin Meng and Lam~For Kwok.
\newblock Adaptive blacklist-based packet filter with a statistic-based
  approach in network intrusion detection.
\newblock {\em Journal of Network \& Computer Applications}, 39(1):83--92,
  2014.

\bibitem{Naderi2015Joza}
Abbas Naderi-Afooshteh, Anh Nguyen-Tuong, Mandana Bagheri-Marzijarani, Jason~D
  Hiser, and Jack~W Davidson.
\newblock Joza: Hybrid taint inference for defeating web application sql
  injection attacks.
\newblock In {\em 2015 45th Annual IEEE/IFIP International Conference on
  Dependable Systems and Networks}, pages 172--183. IEEE, 2015.

\bibitem{Oza2014HTTP}
Aditya Oza, Kevin Ross, Richard~M. Low, and Mark Stamp.
\newblock Http attack detection using n -gram analysis.
\newblock {\em Computers \& Security}, 45(3):242--254, 2014.

\bibitem{Perdisci2009McPAD}
Roberto Perdisci, Davide Ariu, Prahlad Fogla, Giorgio Giacinto, and Wenke Lee.
\newblock Mcpad: A multiple classifier system for accurate payload-based
  anomaly detection.
\newblock {\em Computer Networks the International Journal of Computer \&
  Telecommunications Networking}, 53(6):864--881, 2009.

\bibitem{Robertson2006Using}
William~K. Robertson, Giovanni Vigna, Christopher Kruegel, and Richard~A.
  Kemmerer.
\newblock Using generalization and characterization techniques in the
  anomaly-based detection of web attacks.
\newblock In {\em Network and Distributed System Security Symposium, NDSS 2006,
  San Diego, California, Usa}, 2006.

\bibitem{Song2009Spectrogram}
Yingbo Song, Angelos~D. Keromytis, and Salvatore~J. Stolfo.
\newblock Spectrogram: A mixture-of-markov-chains model for anomaly detection
  in web traffic.
\newblock In {\em Network and Distributed System Security Symposium, NDSS 2009,
  San Diego, California, Usa, February - February}, pages 121--135, 2009.

\bibitem{Swarnkar2016OCPAD}
Mayank Swarnkar and Neminath Hubballi.
\newblock Ocpad: One class naive bayes classifier for payload based anomaly
  detection.
\newblock {\em Expert Systems with Applications}, 64:330--339, 2016.

\bibitem{Wang2006Anagram}
Ke~Wang, Janak~J. Parekh, and Salvatore~J. Stolfo.
\newblock {\em Anagram: a content anomaly detector resistant to mimicry
  attack}.
\newblock Springer Berlin Heidelberg, 2006.

\bibitem{Wang2004Anomalous}
Ke~Wang and Salvatore~J. Stolfo.
\newblock Anomalous payload-based network intrusion detection.
\newblock {\em Lecture Notes in Computer Science}, 3224:203--222, 2004.

\bibitem{Wang2014Autonomic}
Wei Wang, Thomas Guyet, René Quiniou, Marie~Odile Cordier, Florent Masseglia,
  and Xiangliang Zhang.
\newblock Autonomic intrusion detection: Adaptively detecting anomalies over
  unlabeled audit data streams in computer networks.
\newblock {\em Knowledge-Based Systems}, 70:103--117, 2014.

\bibitem{Zhang2015Detecting}
Ji~Zhang, Hongzhou Li, Qigang Gao, Hai Wang, and Yonglong Luo.
\newblock Detecting anomalies from big network traffic data using an adaptive
  detection approach.
\newblock {\em Information Sciences An International Journal}, 318(C):91--110,
  2015.

\bibitem{Zheng2013Path}
Yunhui Zheng and Xiangyu Zhang.
\newblock Path sensitive static analysis of web applications for remote code
  execution vulnerability detection.
\newblock In {\em International Conference on Software Engineering}, pages
  652--661, 2013.

\bibitem{Zhong2015Detecting}
Yang Zhong, Hiroshi Asakura, Hiroki Takakura, and Yoshihito Oshima.
\newblock Detecting malicious inputs of web application parameters using
  character class sequences.
\newblock In {\em Computer Software and Applications Conference (COMPSAC), 2015
  IEEE 39th Annual}, volume~2, pages 525--532. IEEE, 2015.

\end{thebibliography}


\begin{thebibliography}{10}

\bibitem{Ariu2011HMMPayl}
Davide Ariu, Roberto Tronci, and Giorgio Giacinto.
\newblock Hmmpayl : An intrusion detection system based on hidden markov
  models.
\newblock {\em Computers \& Security}, 30(4):221--241, 2011.

\bibitem{Baram2012Online}
Yoram Baram, El~Yaniv Ran, and Kobi Luz.
\newblock Online choice of active learning algorithms.
\newblock {\em Journal of Machine Learning Research}, 5(1):255--291, 2012.

\bibitem{Duda1973Pattern}
By~R~O Duda and P~E Hart.
\newblock {\em Pattern Classification and Scene Analysis, ch}.
\newblock Wiley, 1973.

\bibitem{Perdisci2009McPAD}
Roberto Perdisci, Davide Ariu, Prahlad Fogla, Giorgio Giacinto, and Wenke Lee.
\newblock Mcpad: A multiple classifier system for accurate payload-based
  anomaly detection.
\newblock {\em Computer Networks the International Journal of Computer \&
  Telecommunications Networking}, 53(6):864--881, 2009.

\bibitem{Rendell1987Layered}
Larry Rendell, Raj Sheshu, and David Tcheng.
\newblock Layered concept-learning and dynamically variable bias management.
\newblock In {\em International Joint Conference on Artificial Intelligence},
  pages 308--314, 1987.

\bibitem{rokach2010ensemble}
Lior Rokach.
\newblock Ensemble-based classifiers.
\newblock {\em Artificial Intelligence Review}, 33(1-2):1--39, 2010.

\bibitem{Schohn2000Less}
Greg Schohn and David Cohn.
\newblock Less is more: Active learning with support vector machines.
\newblock In {\em Int Conf}, pages 839--846, 2000.

\bibitem{Sebastiani2002Machine}
Fabrizio Sebastiani.
\newblock Machine learning in automated text categorization.
\newblock {\em Acm Computing Surveys}, 34(1):1--47, 2002.

\bibitem{Tong2001Support}
Simon Tong and Daphne Koller.
\newblock Support vector machine active learning with applications to text
  classification.
\newblock {\em Journal of Machine Learning Research}, 2(1):45--66, 2001.

\bibitem{Vapnik2006Estimation}
Vladimir Vapnik.
\newblock {\em Estimation of Dependences Based on Empirical Data}.
\newblock Springer New York, 2006.

\bibitem{Wolpert1992Stacked}
David~H. Wolpert.
\newblock Stacked generalization *.
\newblock {\em Neural Networks}, 5(2):241--259, 1992.

\end{thebibliography}


\begin{thebibliography}{10}

\bibitem{cavnar1994n}
William~B Cavnar, John~M Trenkle, et~al.
\newblock N-gram-based text categorization.
\newblock {\em Ann Arbor MI}, 48113(2):161--175, 1994.

\bibitem{Fielding1999RFC}
R.~Fielding, Gett~S J., J.~Mogul, H.~Nielsen, L.~Masinter, P.~Leach, and
  T.~Berners-Lee.
\newblock Rfc 2616: Hypertext transfer protocol - http/1.1.
\newblock {\em Computer Science \& Communications Dictionary}, 7(9):3969 --
  3973, 1999.

\bibitem{Jamdagni2013Review}
Aruna Jamdagni, Zhiyuan Tan, Xiangjian He, Priyadarsi Nanda, and Ren~Ping Liu.
\newblock Repids: A multi tier real-time payload-based intrusion detection
  system.
\newblock {\em Computer Networks}, 57(3):811--824, 2013.

\bibitem{Juvonen2015Online}
Antti Juvonen, Tuomo Sipola, and Timo Inen.
\newblock Online anomaly detection using dimensionality reduction techniques
  for http log analysis.
\newblock {\em Computer Networks the International Journal of Computer \&
  Telecommunications Networking}, 91(C):46--56, 2015.

\bibitem{Hillstone}
Hillstone Networks.
\newblock Hillstone e-series next-generation firewalls.
\newblock \url{http://www.hillstonenet.com/}.

\bibitem{Oza2014HTTP}
Aditya Oza, Kevin Ross, Richard~M. Low, and Mark Stamp.
\newblock Http attack detection using n -gram analysis.
\newblock {\em Computers \& Security}, 45(3):242--254, 2014.

\bibitem{Prandl2015A}
Stefan Prandl, Mihai Lazarescu, and Duc~Son Pham.
\newblock {\em A Study of Web Application Firewall Solutions}.
\newblock Springer International Publishing, 2015.

\bibitem{ModSecurity}
Trustwave.
\newblock Modsecurity.
\newblock \url{https://www.modsecurity.org/}.

\bibitem{CRS}
Trustwave.
\newblock Modsecurity core rule set.
\newblock
  \url{https://www.owasp.org/index.php/Category:OWASP_ModSecurity_Core_Rule_Set_Project}.

\bibitem{Wang2006Anagram}
Ke~Wang, Janak~J. Parekh, and Salvatore~J. Stolfo.
\newblock {\em Anagram: a content anomaly detector resistant to mimicry
  attack}.
\newblock Springer Berlin Heidelberg, 2006.

\bibitem{Wang2004Anomalous}
Ke~Wang and Salvatore~J. Stolfo.
\newblock Anomalous payload-based network intrusion detection.
\newblock {\em Lecture Notes in Computer Science}, 3224:203--222, 2004.

\end{thebibliography}


\begin{thebibliography}{10}

\bibitem{Chang2007LIBSVM}
Chih~Chung Chang and Chih~Jen Lin.
\newblock Libsvm: A library for support vector machines.
\newblock {\em Acm Transactions on Intelligent Systems \& Technology}, 2(3,
  article 27):389--396, 2007.

\bibitem{han2011data}
Jiawei Han, Jian Pei, and Micheline Kamber.
\newblock {\em Data mining: concepts and techniques}.
\newblock Elsevier, 2011.

\bibitem{hsu2003practical}
Chih-Wei Hsu, Chih-Chung Chang, Chih-Jen Lin, et~al.
\newblock A practical guide to support vector classification.
\newblock 2003.

\bibitem{Jamdagni2013Review}
Aruna Jamdagni, Zhiyuan Tan, Xiangjian He, Priyadarsi Nanda, and Ren~Ping Liu.
\newblock Repids: A multi tier real-time payload-based intrusion detection
  system.
\newblock {\em Computer Networks}, 57(3):811--824, 2013.

\bibitem{Juvonen2015Online}
Antti Juvonen, Tuomo Sipola, and Timo Inen.
\newblock Online anomaly detection using dimensionality reduction techniques
  for http log analysis.
\newblock {\em Computer Networks the International Journal of Computer \&
  Telecommunications Networking}, 91(C):46--56, 2015.

\bibitem{Kozakevicius2015URL}
Alice Kozakevicius, Cristian Cappo, Bruno~A. Mozzaquatro, Raul~Ceretta Nunes,
  and Christian~E. Schaerer.
\newblock Url query string anomaly sensor designed with the bidimensional haar
  wavelet transform.
\newblock {\em International Journal of Information Security}, 14(6):561--581,
  2015.

\bibitem{Kruegel2003Anomaly}
Christopher Kruegel and Giovanni Vigna.
\newblock Anomaly detection of web-based attacks.
\newblock In {\em ACM Conference on Computer and Communications Security},
  pages 251 -- 261, 2003.

\bibitem{lippmann20001999}
Richard Lippmann, Joshua~W Haines, David~J Fried, Jonathan Korba, and Kumar
  Das.
\newblock The 1999 darpa off-line intrusion detection evaluation.
\newblock {\em Computer networks}, 34(4):579--595, 2000.

\bibitem{Mchugh2000Testing}
John Mchugh.
\newblock Testing intrusion detection systems: a critique of the 1998 and 1999
  darpa intrusion detection system evaluations as performed by lincoln
  laboratory.
\newblock {\em Acm Transactions on Information \& System Security},
  3(4):262--294, 2000.

\bibitem{witten1999weka}
Ian~H Witten, Eibe Frank, Leonard~E Trigg, Mark~A Hall, Geoffrey Holmes, and
  Sally~Jo Cunningham.
\newblock Weka: Practical machine learning tools and techniques with java
  implementations.
\newblock 1999.

\end{thebibliography}


\begin{thebibliography}{1}

\bibitem{Maggi2009Protecting}
Federico Maggi, William Robertson, Christopher Kruegel, and Giovanni Vigna.
\newblock Protecting a moving target: Addressing web application concept drift.
\newblock In {\em International Symposium on Recent Advances in Intrusion
  Detection}, pages 21--40, 2009.

\bibitem{Schlimmer1986Beyond}
Jeffrey~C. Schlimmer and Richard~H. Granger.
\newblock Beyond incremental processing: Tracking concept drift.
\newblock In {\em National Conference on Artificial Intelligence. Philadelphia,
  Pa, August 11-15, 1986. Volume 1: Science}, pages 502--507, 1986.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{url}
\usepackage[hidelinks]{hyperref}
\usepackage{amsmath}
\usepackage{multirow}
\usepackage{hhline}
\usepackage{graphicx}
\usepackage{float}
\usepackage{caption}
\usepackage{subfigure}
\usepackage{textcomp}
\usepackage{dblfloatfix}

\begin{document}

\section{Introduction}\label{introduction}

Web attacks are attacks that exclusively use the HTTP/HTTPS protocol.
Symantec Internet Security Threat Report (ISTR) \cite{ISTR2016} reveals that the number of web attacks explosively increased, reaching 1.1 million every day in 2015, more than twice the rate in 2014 (0.493 million).
Among web attacks, code injection attacks against web applications \cite{Fonseca2014Evaluation} increased each year and accounted for at least 96.15\% of web attacks in 2015, according to the Imperva Web Application Attack Report (WAAR) \cite{WAAR2014,WAAR2015,WAAR2016}.
In 2015, Team GhostShell claimed to have hacked numerous websites using SQL injection (SQLI) attacks \cite{lawal2016systematic}, and disclosed thousands of compromised account details, including emails, user names, addresses, telephone numbers, and other privacy information \cite{Ghost}.

This paper focuses on the most frequent types of web-based code-injection attacks in 2015 \cite{WAAR2016}, namely, Cross-Site Scripting (XSS) attack \cite{hydara2015current}, SQLI attack, Directory Traversal (DT) \cite{WAAR2016}, and Remote File Inclusion (RFI) \cite{WAAR2016}, respectively accounting for 49.09\%, 28.32\%, 9.82\% and 8.92\% of the entire web attacks.
Attack vectors of these attacks exist in user input.
Whenever user input is improperly handled, these attacks can succeed.
Since most user input data exists in queries \footnote{Web request query string is consistently termed as query in this paper for simplicity.}, detecting malicious queries is the core of detecting web-based code injection attacks.
For web applications whose source code is unavailable or difficult to obtain, Web Intrusion Detection Systems (Web IDSs) are the only option \cite{Prokhorenko2016Web}.
A web IDS acts as an intermediate layer between protected web applications and users, and analyzes web traffic to detect possibly malicious activities.

A considerable amount of effort has been made to detect malicious queries in web requests using web IDSs.
Two major detection approaches are utilized: signature-based detection and anomaly-based detection.
A popular signature-based detection approach is WAF, which maintains a signature library of known signatures, and compares new web requests with the signatures.
Signature-based approaches are efficient in detecting known attacks with a low False Positive (FP) rate, but unable to detect previously unknown attacks(e.g., zero-day attacks).
Anomaly-based detection approaches \cite{Kruegel2003Anomaly,Kruegel2005A,Robertson2006Using,Song2009Spectrogram,Kozakevicius2015URL,Juvonen2015Online} usually rely on a detection model to identify anomalous web requests.
In contrast to signature-based approaches, anomaly-based approaches can detect unknown attacks, but with a high FP rate.
The two approaches are often applied in a complementary fashion in web IDSs: anomaly detection serves to discover unknown attacks, whose attack signatures are then utilized by signature-based approach for detection.
The ability to detect unknown attacks has drawn wide academic attention to anomaly-based approaches.

However, existing anomaly-based approaches for web attack detection are constant, in other words, they usually collect all training data once to induce a constant detection model.
Since attackers have become more sophisticated and utilized more advanced web attack toolkits, the detection model might become obsolete and outdated, incapable of detecting the latest malicious queries in web attacks.
Previous adaptive attack detection methods \cite{Meng2014Adaptive,Wang2014Autonomic,Zhang2015Detecting} are designed for network intrusions and are not applicable to web attacks.
A practical solution for keeping the web attack detection model constantly updated is to incorporate the latest important queries, including informative benign queries and representative malicious queries.
The existing approach to obtaining the latest important queries is randomly selecting requests from web traffic and then manually labeling them.
Unfortunately, the majority of randomly selected queries are similar benign ones, and there is low probability of a query being important.
Since web traffic is huge, it does take considerable manual labeling work to obtain important queries.
This challenge has motivated our research work to build a malicious query detection system that can adaptively incorporate the latest important queries to update the detection model.

In this paper, we address the issues related to the adaptive detection of malicious queries in web attacks.
We present AMODS, an adaptive system for this purpose.
AMODS leverages an efficient adaptive learning strategy, SVM HYBRID, which is a hybrid of Suspicion Selection (SS) and Exemplar Selection (ES).
The former features in acquiring the most important informative queries, namely, suspicions, while the latter specializes in harvesting representative malicious queries, exemplars.
Suspicions and exemplars are called important queries.
AMODS aims to identify attacks as early as possible by periodically checking the latest important queries.
The number of important queries is trivial, so manual labeling work is reduced to the minimum.
The important queries are then incorporated into the training pool to update the detection model, which is a stacking-based ensemble classifier, composed of three base classifiers and a meta classifier.
SVM \cite{Vapnik2006Estimation} is used as its meta classifier, so that SVM HYBRID can exert the detection model to obtain important queries.
A ten-day query dataset collected from an academic institute website's web server logs is used for evaluation.
Since the logs record requests passing through a commercial WAF, malicious queries obtained by our system can be used to update the WAF signature library.

The main contributions of this study are as follows:

\begin{itemize}
\item We present AMODS, an adaptive system for detecting malicious queries in web attacks. To the best of our knowledge, this is the first work in the adaptive detection of web attacks.

\item We propose SVM HYBRID, an adaptive learning strategy designed for the efficient selection of important queries.

\item We employ stacking, a meta-learning \cite{brazdil2008metalearning} based ensemble classifier, as the underlying detection model for accurate malicious query detection.

\item AMODS outperforms existing web attack detection methods with an F-value of 94.79\% and FP-rate of 0.09\% on a real-world ten-day query set. The total number of malicious queries obtained by SVM HYBRID is 2.78 times that by SVM AL during the ten-day detection using AMODS.
\end{itemize}

The remainder of the paper is organized as follows.
Section \ref{Section2.backRela} provides background and related work.
Detailed design and implementation of the system, together with the proposed adaptive learning strategy, namely, SVM HYBRID, will be presented in Section \ref{Section3.design}.
Section 4 shows data collection and preprocessing.
Next, Section \ref{Section4.experiment} introduces the process of the determination of feature reduction and stacked classifiers, then shows the efficacy of our system.
In addition, AMODS is compared with constant models and adaptive models, including SVM AL and random selection, in terms of detection performance and the number of malicious queries obtained.
Detection performance comparisons with related work are also demonstrated.
Afterwards, we clarify some relevant issues of our approach in Section \ref{Section5.discussion}.
Finally, Section \ref{Section6.conclusion} concludes this paper.

\bibliographystyle{plain}

\bibliography{mybibfile}

\end{document}